# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 6.3 Wall-crossing
The number $W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}}(t))$ remains constant in each of the wall-crossing events described in Lemma 6.2(2).
Proof. All local considerations in the proof of Theorems 3.2 and 3.7 can literally be transferred to our situation. Thus, we only confirm the global transversality relations which all take form of certain cohomology vanishing conditions.
(1) The wall-crossing described in Lemma 6.2(2i) coincides with that in Lemma 5.2(2ii). The transversality condition in the treatment of this wall-crossing reads (cf. formula (23))
$H^{0}(S_{2},{\mathcal{O}}_{S_{2}}({\boldsymbol{d}}-p_{1}-p_{2}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}(S_{2}),$ (28)
@@
Thus, the Riemann-Roch formula yields $h^{0}(S_{2},{\mathcal{O}}_{S_{2}}({\boldsymbol{d}}))=2$, and hence (28) holds due to the general position of $w_{1},w_{2}$ on $C={\Phi}(S_{2})$.
(2) The wall-crossing described in Lemma 6.2(2ii) coincides with that in Lemma 5.2(2iii) (the case of $C$ unibranch at $w_{i}^{\sigma}(t^{*})=w_{j}^{\sigma}(t^{*})$, i.e., $p_{i}^{\sigma}=p_{j}^{\sigma}$). The required transversality condition coincides with relation (28) confirmed in the preceding paragraph.
(3) The wall-crossing described in Lemma 6.2(2iii) coincides with that in Lemma 5.2(2iv), the subcase of ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(\Delta)$, $i=1,2$, birational immersions. The local study is a word-for-word copy of the proof of Lemma 4.3, part (5a) (restricted to the case of birational immersions). Similarly to the proof of Lemma 4.3, part (5b), and the proof of Lemma 5.3, part (4), the global transversality condition amounts to
@@
To verify the above $h^{0}$-vanishing conditions we argue as in step (2) above.
(4) The last wall-crossing to consider is associated with the wall described in Lemma 6.2(2v). That wall is very similar to the wall appearing in Lemma 5.2(2v)]. The local study of the wall-crossing is a word-for-word copy of the study performed in the proof of Lemma 5.3, part (5). Thus, we only address the global transversality condition analogous to that in formula (26):
$h^{0}({\mathbb{E}},{\mathcal{N}}_{{\Phi}^{\prime}}({\boldsymbol{d}}))=0,$
@@
where ${\Phi}^{\prime}:{\mathbb{E}}\to\widehat{C}\overset{{\Phi}}{\to}{\mathbb{P}}^{2}$ is ${\Phi}$ combined with the normalization, ${\mathcal{N}}_{{\Phi}^{\prime}}$ is the normal bundle, $p_{i^{\prime}}^{(s^{\prime})},p_{i^{\prime\prime}}^{(s^{\prime\prime})}$ are mobile points in $\widehat{\boldsymbol{w}}_{\partial}$ (while the other points are fixed), and $p^{\prime},p^{\prime\prime}$ are the preimages of the node $p$ in ${\mathbb{E}}$. It is easy to see that
$\deg{\boldsymbol{d}}=0.$
However, the points $w_{1},w_{2}$ are in general position on $C={\Phi}(\widehat{C})$, and hence $p_{1},p_{2}$ are in general position on ${\mathbb{E}}$, which by the Riemann-Roch formula implies the desired $h^{0}$-vanishing. $\Box$

### v2

@@ 6.3 Wall-crossing
The numbers $W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}}(t))$ remain constant in each of the wall-crossing events described in Lemma 6.2(2).
Proof. All local considerations in the proof of Theorems 3.2 and 3.6 can literally be transferred to our situation. Thus, we only confirm the global transversality relations which all take form of certain cohomology vanishing conditions.
(1) The wall-crossing described in Lemma 6.2(2ii) coincides with that in Lemma 5.2(2ii). The transversality condition in the treatment of this wall-crossing reads (cf. formula (36))
$H^{0}(\widehat{C},{\mathcal{O}}_{\widehat{C}}({\boldsymbol{d}}-p_{1}-p_{2}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}(\widehat{C}),$ (45)
@@
Thus, the Riemann-Roch formula yields $h^{0}(\widehat{C},{\mathcal{O}}_{\widehat{C}}({\boldsymbol{d}}))=2$, and hence (45) holds due to the general position of $w_{1},w_{2}$ on $C={\Phi}(\widehat{C})$.
(2) The wall-crossing described in Lemma 6.2(2iii) coincides with that in the proof of Lemma 5.2(2iii). The required transversality condition coincides with relation (45) confirmed in the preceding paragraph.
(3) The wall-crossing described in Lemma 6.2(2iv) coincides with that in Lemma 5.2(2iv), the subcase of ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(\Delta)$, $i=1,2$, birational immersions. The local study is a word-for-word copy of the proof of Lemma 4.3, part (5a) (restricted to the case of birational immersions). Similarly to the proof of Lemma 4.3, part (5b), and the proof of Lemma 5.3, part (4), the global transversality condition amounts to
@@
Thus, the required $h^{1}$-vanishing holds.
(4) The last wall-crossing to consider is associated with the wall described in Lemma 6.2(2v). That wall is very similar to the wall appearing in Lemma 5.2(2v). The local study of the wall-crossing is a word-for-word copy of the study performed in the proof of Lemma 5.3, part (5). Thus, we only address the global transversality condition analogous to that in formula (40):
$h^{0}(\widehat{C},{\mathcal{N}}_{{\Phi}^{\prime}}({\boldsymbol{d}}))=0,$
@@
where ${\Phi}^{\prime}:\widehat{C}\to\widehat{C}\overset{{\Phi}}{\to}{\mathbb{P}}^{2}$ is a composition of ${\Phi}$ with the normalization, ${\mathcal{N}}_{{\Phi}^{\prime}}$ is the normal bundle, $p_{i^{\prime}}^{(s^{\prime})},p_{i^{\prime\prime}}^{(s^{\prime\prime})}$ are mobile points in $\widehat{\boldsymbol{w}}_{\partial}$ (while the other points are fixed), and $p^{\prime},p^{\prime\prime}$ are the preimages of the node $p$ in $\widehat{C}$. It is easy to see that
$\deg{\mathcal{N}}_{{\Phi}^{\prime}}({\boldsymbol{d}})=0.$
However, the points $w_{1},w_{2}$ are in general position on $C={\Phi}(\widehat{C})$, and hence $p_{1},p_{2}$ are in general position on $\widehat{C}$, which by the Riemann-Roch formula implies the desired $h^{0}$-vanishing. $\Box$